OpenStudy (anonymous):

MEDAL AND FAN What is the equation of the line that passes through the points (-3, 2) and (-5, 8)? A) 3x-y=-11 B) 3x+y=-7 C) 3x+y=5 D) x+3y=3

OpenStudy (anonymous):

@misty1212 could you please help me with this one as well?

OpenStudy (misty1212):

yes, slope we should find first

OpenStudy (anonymous):

y-mx+b

OpenStudy (anonymous):

-3?

OpenStudy (misty1212):

we do \[\frac{8-2}{-5-(-3)}\] to find the slope

OpenStudy (anonymous):

8-2/-5+3 which would be 6/-2 =3 correct?

OpenStudy (anonymous):

-3 excuse me

OpenStudy (misty1212):

it would be \(\frac{6}{-2}=-3\) yes

OpenStudy (anonymous):

Alright what do we figure out next?

OpenStudy (misty1212):

we can do \[y-y_1=m(x-x_1)\] with \(m=-3, x_1=-3,y_1=2\)

OpenStudy (misty1212):

\[y-2=-3(x+3)\] if we substitute then multiply \[y-2=-3x-9\]

OpenStudy (anonymous):

add 2 to both sides?

OpenStudy (misty1212):

yeah, also \(3x\) to both sides

OpenStudy (anonymous):

So 3x+y=-11 :3

OpenStudy (anonymous):

no thats not an answer choice....

OpenStudy (misty1212):

\(-9+2=-7\)

OpenStudy (anonymous):

Oh! :D so it would be y+3x=-7 answer choice B :D

OpenStudy (anonymous):

Thank you so much @misty1212 ! :3

OpenStudy (misty1212):

happy